In C. elegans, a recessive mutation that causes the worms to twitch is 6 map units from a recessive mutation that results in a dumpy phenotype. A twitcher dumpy hermaphrodite is mated to a wild-type male. A wild-type F1 male is found and test-crossed to the twitcher dumpy hermaphrodite.

a) What fraction of the offspring is predicted to be twitcher, but not dumpy, males?
Answer #1

T = Dominant allele = Non-twitch
t = Recessive allele = Twitch

D = Dominant allele = Normal
d = Recessive allele = Dumpy

Parental cross: ttdd X TTDD
F1 progeny: TtDd
F1 test cross: TtDd X ttdd
Distance between T and D = 6 mu
i.e. Recombinationfrequency = 6%
Recombinants = Td/td + tD/td = 6%
Td/td = tD/td = 3%

Parental combinations = TD/td + td/td = 94%
TD/td = td/td = 47%

Freequency of twitcher, non-dumpy = tD/td = 3%
Probability for a male = 1/2 = 0.5
Overall probability = 3*0.5 = 1.5%